# Causal Inference: A Tale of Three Frameworks

### v1

@@ 3.3 Potential Outcomes and DAGs: Single World Intervention Graphs (SWIGs)
The potential outcomes and DAG frameworks can be unified through single world intervention graphs (SWIGs) (Richardson and Robins,, 2013). SWIGs explicitly encode potential outcomes on the graph by splitting nodes into fixed (intervened) and random components. This representation enables two key features: (1) conditional independence statements involving both observed and potential outcomes can be read using a modified d-separation rule, and (2) identification results equivalent to do-calculus can be derived from these conditional independence conditions via a simple reformulation known as the potential outcome (po) calculus (Malinsky et al.,, 2019; Shpitser et al.,, 2022). This reformulation shows that from the perspective of potential outcomes, the po-calculus (and hence the do-calculus) is reducible to conditional independence, recursive substitution and causes only influence descendants in the SWIG. For completeness, we note that SWIGs can also be defined via a local Markov property with respect to a total ordering of the variables, without explicit reference to a graphical representation (see Richardson and Robins,, 2023). The same potential outcome model was originally termed the Finest Fully Randomized Causally Interpretable Structured Tree Graph (FFRCISTG); see Robins, (1986) and Richardson and Robins, (2013, Appendix C); note that the “tree graphs” referred to here are “event trees,” not DAGs.
@@
Pearl, (2009, Ex. 11.3.3) claimed that the causal DAG model in Figure 4(a) does not imply (17). However, as can be seen from the SWIG in Figure 4(b), the relationship $Y(a,b)\perp\!\!\!\perp B(a)\mid Z(a),A=a$ holds, which implies (17) through the consistency assumption.
@@ 4 Comparison between the three frameworks
Although, as we have shown, these frameworks often lead to overlapping mathematical results, they arise from different philosophical traditions and emphasize different aspects of the causal enterprise (Pearl,, 2009, § 1.4). The potential outcomes framework adopts a minimalist or design-based view of causality, focusing on comparisons between hypothetical outcomes under different treatment conditions, such as Y(1)Y(0)Y(1)-Y(0). This approach does not require specifying how all outcomes are generated from other variables in the system. It is particularly attractive in settings where causal identification relies on features of the study design, such as randomization or the use of instrumental variables, rather than on a full model of the data-generating process. Because of its minimal assumptions and broad applicability, this framework has been described by some authors as the most general approach to causal inference (Imbens and Rubin,, 2015; Rosenbaum,, 2010).
@@
In contrast, NPSEM-IEs adopt a more mechanistic perspective, rooted in classical physics and engineering, where causality is represented through deterministic functional relationships between variables and exogenous noise terms. In this tradition, causal systems are fully specified via structural equations, and randomness reflects ignorance about latent factors rather than intrinsic stochasticity. This view aligns with Laplace’s deterministic vision of natural laws (Laplace,, 1825) and provides a rich language for reasoning about complex causal systems.
@@
Causal DAGs provide a bridge between this minimalist perspective and mechanistic modeling. By encoding assumptions about independencies, DAGs help justify potential-outcomes-based identifiability conditions (e.g., ignorability) while avoiding over-specification of functional forms. For instance, a DAG makes it transparent why randomization implies that no arrows point into $A$. This, in turn, ensures the independence $A\perp\!\!\!\perp Y(a),\,a\in\{0,1\}$, which can be read off from a SWIG intervening on $A$. Moreover, DAGs clarify backdoor adjustment sets for observational studies (Pearl,, 2009). In contrast to NPSEMs, all relationships in causal DAG models are formulated as inherently stochastic. As Pearl, (2009, p. 26) observed, this reflects the modern conception of physics, sometimes described as quantum mechanical, in which probabilistic laws replace strict determinism as a more realistic description of physical processes. At the same time, DAGs retain the ability to articulate complex structural relationships, making them a powerful tool for both identification and hypothesis generation.
@@
As a caveat, we note that although Pearl’s analogy draws on the idea that modern physics views natural laws as fundamentally probabilistic rather than strictly deterministic, it should be interpreted only metaphorically: causal DAGs are classical probabilistic models that admit hidden variable representations, whereas quantum systems in general do not. Quantum phenomena require richer nested probabilistic frameworks, sometimes viewed as supermodels of the classical theory, that extend beyond the scope of standard DAGs (Richardson et al.,, 2023).
@@
4.2 Single-world versus cross-world causal inference
@@
The key difference between the causal DAG framework with its associated “do” notation and the other two frameworks (NPSEM-IE and potential outcomes) is that the former does not provide a way to directly express counterfactual quantities that condition on observed variables rather than interventions, such as $P(Y(a)\mid A=a^{\prime})$ (cf. Pearl,, 2009, §1.4.4). An important example is the ETT, $\mathbb{E}[Y(1)-Y(0)\mid A=1]$, which is not expressible in the “do” notation. Compared with the causal DAG model, the NPSEM-IE model also imposes many more assumptions due to the additional cross-world assumptions. We now elaborate this point.
@@
Because the independent error assumptions in NPSEM, when taken at face value, resemble the independent error assumptions commonly made in statistical modeling, they are often perceived as mild by statisticians who are not familiar with the causal modeling framework. Interestingly, many of the same statisticians would view the assumption of no unmeasured confounding as an extremely strong assumption. Before the recent evolution in causal inference, it was standard practice in statistics to avoid causal claims altogether, precisely because of the concern that confounding renders correlation insufficient for establishing causation (Pearl et al.,, 2016, Preface).
@@
where $\mathcal{L}$ denotes the support of $L$ (Pearl,, 2009, p. 101 footnote 14). Basic algebra then yields the strong ignorability condition,
@@
In fact, we have the following general result due to Pearl, (2009, p. 101, eqn. (3.56) & footnote 14).
@@
To understand the number of extra assumptions imposed by the NPSEM-IE, consider again the NPSEM-IE in eqn. (4) and a simple case where $L$, AA, and $Y$ are all binary. A full model for the joint distribution of $(L,A(l),Y(l,a))$, for $a\in\{0,1\}$ and $l\in\{0,1\}$, has dimension $2^{1+2+4}-1=127$. The no-unmeasured-confounding assumption in eqn. (1) reduces this to a 123-dimensional model. In contrast, the strong ignorability assumption in eqn. (19) further reduces the model to 121 dimensions, and the independent error assumption of the NPSEM-IE implies a model of only 19 dimensions! In other words, NPSEM-IE imposes an additional 12319=104123-19=104 constraints, all of which are not necessary for the identification of the ACE (Robins and Richardson,, 2011). Most of these constraints, such as $A(l=1)\perp\!\!\!\perp Y(l=0,a)$, are examples of cross-world assumptions, which remain untestable even under randomized treatment assignment. These assumptions violate the principle of “no causation without manipulation” (Holland,, 1986) and undermine the connection between experimentation and causal inference (Robins,, 2003; Robins and Richardson,, 2011).
@@
In some cases, the additional independence assumptions in NPSEM-IE lead to new identification results. For example, Figure 5 shows the mediation setting with $A$ the treatment, $M$ the mediator, and $Y$ the outcome of interest. An estimand of interest in this setting is known as the natural (or pure) direct effect (Robins and Greenland,, 1992), defined as NDE=E[Y(1,M(0))Y(0)]NDE=E[Y(1,M(0))-Y(0)]. The NDE can be identified under the NPSEM-IE using the mediation formula (Pearl,, 2001)
@@
Other examples include the identification of the local average treatment effect in instrumental variable models (Angrist et al.,, 1996), tighter bounds on natural indirect effects in mediation analysis without assuming cross-world independencies (Imai et al.,, 2010), and the identification of, or tighter bounds on, the survivor average treatment effect (Ding et al.,, 2011; Wang et al., 2017a, ; Wang et al., 2017b, ).
@@
    (Linear, non-Gaussian) Shimizu et al., (2006) showed that if the functions $f_{j}$ in the NPSEM-IE (20) are linear in $U_{j}$ and $\varepsilon_{j}$ with nonzero coefficients, and if the error terms $\varepsilon_{j}$ are non-Gaussian with strictly positive density, then the underlying graph structure is identifiable from the observed data distribution. This model is known as the linear non-Gaussian acyclic model (LiNGAM).
@@
    Hoyer et al., (2008) showed that, except in “rare” cases, a distribution does not admit an additive noise model structure in both directions simultaneously; see also Peters et al., (2017, Theorem 4.5). In fact, if the noise $\varepsilon_{Y}$ is Gaussian, then only linear functions allow for additive noise models in both directions (Hoyer et al.,, 2008, Corollary 1). More broadly, Peters et al., (2014) showed that if the functions $f_{j}$ in the NPSEM-IE (3) are three times differentiable and nonlinear in any component of $U_{j}$ (see Peters et al.,, 2014, Theorem 7.7 for a precise definition), then the corresponding graph structure is identifiable from the observed data distribution. These results extend to the so-called post-nonlinear models of the form $V_{j}=g_{j}(f_{j}(U_{j})+\varepsilon_{j})$. These models also generally exist in at most one direction, except in rare cases (Zhang and Hyvärinen,, 2009).
Identification of the causal structure in linear models with Gaussian additive noise is more challenging. For example, consider again the bivariate causal model in (21). If $f_{Y}(X)$ is linear in $X$, then there exist $\beta$ and $\varepsilon_{X}$ such that $X=\beta Y+\varepsilon_{X}$ and $Y\perp\!\!\!\perp\varepsilon_{X}$ if and only if both $\varepsilon_{Y}$ and $X$ are Gaussian (Peters et al.,, 2017, Theorem 4.2). More generally, identification of the causal graph under linear models with Gaussian additive noise is possible if the error terms $\varepsilon_{j}$ are i.i.d., that is, the noise variances are constant across $j$ (Peters and Bühlmann,, 2014). However, the equal variance assumption is not robust to rescaling of the variables.
We conclude this example with a caveat: in the nonlinear case, additive noise models are generally not closed under marginalization, even when marginalizing over unobserved variables. For instance, if $Y=f_{Y}(X)+\varepsilon_{Y}$ and $Z=f_{Z}(Y)+\varepsilon_{Z}$, then the marginal distribution $P_{X,Z}$ does not necessarily admit an additive noise model from $X$ to $Z$. This limitation restricts the applicability of additive noise models, as in most, if not all, scenarios there are intermediate variables on a causal path that are unobserved (Peters et al.,, 2017, §7.1.2). Therefore, additive models should perhaps be viewed as approximations to the true data-generating process rather than exact representations.
@@ 5 Summary and practical recommendations
The NPSEM framework formalizes causal systems using deterministic functional relationships and exogenous error terms. This allows for a generative view of how outcomes arise from their causes, enabling researchers to articulate and evaluate identification assumptions in terms of underlying data-generating mechanisms. As emphasized by Pearl, (2009, p. 244), structural models and their associated graphs are particularly useful as tools for expressing assumptions about cause-effect relationships, as they provide a natural and intuitive representation of mechanistic knowledge. Counterfactual independence assumptions that may appear opaque or difficult to interpret when stated in the potential outcomes notation, such as the ignorability assumption, can often be given immediate and process-based interpretations when expressed structurally. The NPSEM framework also facilitates encoding assumptions on the functional form or even parameter values of the structural functions. However, when paired with the assumption of independent errors (NPSEM-IE), this framework often imposes strong and untestable cross-world assumptions that go beyond what is needed for identification and remain untestable even under randomized treatment assignment. These added assumptions can constrain the space of admissible models and lead to overly rigid inferences.
@@
In this workflow, structural assumptions are naturally expressed in graphical or NPSEM terms, then translated into potential outcomes notation, and finally subjected to algebraic derivation (Pearl,, 2009, p. 245) to isolate those that are strictly necessary for identification. In this way, NPSEMs and their associated graphs (see Section 3.2 for details) serve as tools for evaluating the plausibility or implausibility of conditional independence assumptions stated in terms of potential outcomes. Contemporary implementations such as the target trial emulation framework (Hernán et al.,, 2022) illustrate how structured causal roadmaps can help prevent common inferential pitfalls in observational studies. We refer readers to Wang et al., 2017b ; Wang and Tchetgen Tchetgen, (2018); Yang et al., (2019); Zhou et al., (2024); Dong et al., (2025) for examples that follow this workflow.
We now illustrate these steps in detail using Wang and Tchetgen Tchetgen, (2018). Their study begins by defining the causal estimand, the ACE, in potential outcomes notation, thereby clarifying the scientific target of inference (Step 1). They then express structural assumptions through a DAG and a corresponding SWIG with bidirected arrows, explicitly distinguishing identification assumptions such as the exclusion restriction and independence from modeling assumptions (Step 2). Using algebraic derivations, they establish identification of the ACE under new no-interaction assumptions (A5.a or A5.b), showing that the estimand can be represented as the average Wald estimand (Step 3). Finally, they develop a series of bounded, efficient, and multiply robust estimators that remain consistent if any one of several working models is correctly specified (Step 4). This example demonstrates how potential outcomes, graphical models, and semiparametric theory can be systematically combined into a coherent causal workflow.
Recent advances continue to expand the reach of causal inference beyond traditional domains. Active research areas include causal discovery algorithms for learning causal structure from data (e.g., Spirtes et al.,, 2000; Peters et al.,, 2017), fairness in artificial intelligence (e.g., Kusner et al.,, 2017), and methods for robust learning and generalization across heterogeneous environments (e.g., Peters et al.,, 2016). Related ideas have also influenced reinforcement learning and decision-making in artificial intelligence, where causal reasoning is increasingly recognized as essential for interpretability and transferability. A full treatment of these topics is beyond the scope of this review, but they exemplify how causal thinking continues to shape contemporary research at the intersection of statistics, computer science, and the social and health sciences.

### v2

@@ 3.3 Potential Outcomes and DAGs: Single World Intervention Graphs (SWIGs)
The potential outcomes and DAG frameworks can be unified through single world intervention graphs (SWIGs) (Richardson and Robins, 2013). SWIGs explicitly encode potential outcomes on the graph by splitting nodes into fixed (intervened) and random components. This representation enables two key features: (1) conditional independence statements involving both observed and potential outcomes can be read using a modified d-separation rule, and (2) identification results equivalent to do-calculus can be derived from these conditional independence conditions via a simple reformulation known as the potential outcome (po) calculus (Malinsky et al., 2019; Shpitser et al., 2022). This reformulation shows that from the perspective of potential outcomes, the po-calculus (and hence the do-calculus) is reducible to conditional independence, recursive substitution and causes only influence descendants in the SWIG. For completeness, we note that SWIGs can also be defined via a local Markov property with respect to a total ordering of the variables, without explicit reference to a graphical representation (see Richardson and Robins, 2023). The same potential outcome model was originally termed the Finest Fully Randomized Causally Interpretable Structured Tree Graph (FFRCISTG); see Robins (1986) and Richardson and Robins (2013, Appendix C); note that the “tree graphs” referred to here are “event trees,” not DAGs.
@@
Pearl (2009, Ex. 11.3.3) claimed that the causal DAG model in Figure 4(a) does not imply (17). However, as can be seen from the SWIG in Figure 4(b), the relationship $Y(a,b)\perp\!\!\!\perp B(a)\mid Z(a),A=a$ holds, which implies (17) through the consistency assumption.
@@ 4 Comparison Between the Three Frameworks
Although, as we have shown, these frameworks often lead to overlapping mathematical results, they arise from different philosophical traditions and emphasize different aspects of the causal enterprise (Pearl, 2009, § 1.4). The potential outcomes framework adopts a minimalist or design-based view of causality, focusing on comparisons between hypothetical outcomes under different treatment conditions, such as Y(1)Y(0)Y(1)-Y(0). This approach does not require specifying how all outcomes are generated from other variables in the system. It is particularly attractive in settings where causal identification relies on features of the study design, such as randomization or the use of instrumental variables, rather than on a full model of the data-generating process. Because of its minimal assumptions and broad applicability, this framework has been described by some authors as the most general approach to causal inference (Imbens and Rubin, 2015; Rosenbaum, 2010).
@@
In contrast, NPSEM-IEs adopt a more mechanistic perspective, rooted in classical physics and engineering, where causality is represented through deterministic functional relationships between variables and exogenous noise terms. In this tradition, causal systems are fully specified via structural equations, and randomness reflects ignorance about latent factors rather than intrinsic stochasticity. This view aligns with Laplace’s deterministic vision of natural laws (Laplace, 1825) and provides a rich language for reasoning about complex causal systems.
@@
Causal DAGs provide a bridge between this minimalist perspective and mechanistic modeling. By encoding assumptions about independencies, DAGs help justify potential-outcomes-based identifiability conditions (e.g., ignorability) while avoiding over-specification of functional forms. For instance, a DAG makes it transparent why randomization implies that no arrows point into $A$. This, in turn, ensures the independence $A\perp\!\!\!\perp Y(a),\,a\in\{0,1\}$, which can be read off from a SWIG intervening on $A$. Moreover, DAGs clarify backdoor adjustment sets for observational studies (Pearl, 2009). In contrast to NPSEMs, all relationships in causal DAG models are formulated as inherently stochastic. As Pearl (2009, p. 26) observed, this reflects the modern conception of physics, sometimes described as quantum mechanical, in which probabilistic laws replace strict determinism as a more realistic description of physical processes. At the same time, DAGs retain the ability to articulate complex structural relationships, making them a powerful tool for both identification and hypothesis generation.
@@
As a caveat, we note that although Pearl’s analogy draws on the idea that modern physics views natural laws as fundamentally probabilistic rather than strictly deterministic, it should be interpreted only metaphorically: causal DAGs are classical probabilistic models that admit hidden variable representations, whereas quantum systems in general do not. Quantum phenomena require richer nested probabilistic frameworks, sometimes viewed as supermodels of the classical theory, that extend beyond the scope of standard DAGs (Richardson et al., 2023).
@@
4.2 Single-World vs. Cross-World Causal Inference
@@
The key difference between the causal DAG framework with its associated “do” notation and the other two frameworks (NPSEM-IE and potential outcomes) is that the former does not provide a way to directly express counterfactual quantities that condition on observed variables rather than interventions, such as $P(Y(a)\mid A=a^{\prime})$ (cf. Pearl, 2009, §1.4.4). An important example is the ETT, $\mathbb{E}[Y(1)-Y(0)\mid A=1]$, which is not expressible in the “do” notation. Compared with the causal DAG model, the NPSEM-IE model also imposes many more assumptions due to the additional cross-world assumptions. We now elaborate this point.
@@
Because the independent error assumptions in NPSEM, when taken at face value, resemble the independent error assumptions commonly made in statistical modeling, they are often perceived as mild by statisticians who are not familiar with the causal modeling framework. Interestingly, many of the same statisticians would view the assumption of no unmeasured confounding as an extremely strong assumption. Before the recent evolution in causal inference, it was standard practice in statistics to avoid causal claims altogether, precisely because of the concern that confounding renders correlation insufficient for establishing causation (Pearl et al., 2016, Preface).
@@
where $\mathcal{L}$ denotes the support of $L$ (Pearl, 2009, p. 101 footnote 14). Basic algebra then yields the strong ignorability condition,
@@
In fact, we have the following general result due to Pearl (2009, p. 101, eqn. (3.56) & footnote 14).
@@
To understand the number of extra assumptions imposed by the NPSEM-IE, consider again the NPSEM-IE in eqn. (4) and a simple case where $L$, AA, and $Y$ are all binary. A full model for the joint distribution of $(L,A(l),Y(l,a))$, for $a\in\{0,1\}$ and $l\in\{0,1\}$, has dimension $2^{1+2+4}-1=127$. The no-unmeasured-confounding assumption in eqn. (1) reduces this to a 123-dimensional model. In contrast, the strong ignorability assumption in eqn. (19) further reduces the model to 121 dimensions, and the independent error assumption of the NPSEM-IE implies a model of only 19 dimensions! In other words, NPSEM-IE imposes an additional 12319=104123-19=104 constraints, all of which are not necessary for the identification of the ACE (Robins and Richardson, 2011). Most of these constraints, such as $A(l=1)\perp\!\!\!\perp Y(l=0,a)$, are examples of cross-world assumptions, which remain untestable even under randomized treatment assignment. These assumptions violate the principle of “no causation without manipulation” (Holland, 1986) and undermine the connection between experimentation and causal inference (Robins, 2003; Robins and Richardson, 2011).
@@
In some cases, the additional independence assumptions in NPSEM-IE lead to new identification results. For example, Figure 5 shows the mediation setting with $A$ the treatment, $M$ the mediator, and $Y$ the outcome of interest. An estimand of interest in this setting is known as the natural (or pure) direct effect (Robins and Greenland, 1992), defined as NDE=E[Y(1,M(0))Y(0)]NDE=E[Y(1,M(0))-Y(0)]. The NDE can be identified under the NPSEM-IE using the mediation formula (Pearl, 2001)
@@
Other examples include the identification of the local average treatment effect in instrumental variable models (Angrist et al., 1996), tighter bounds on natural indirect effects in mediation analysis without assuming cross-world independencies (Imai et al., 2010), and the identification of, or tighter bounds on, the survivor average treatment effect (Ding et al., 2011; Wang et al., 2017a, b).
@@
    (Linear, non-Gaussian) Shimizu et al. (2006) showed that if the functions $f_{j}$ in the NPSEM-IE (20) are linear in $U_{j}$ and $\varepsilon_{j}$ with nonzero coefficients, and if the error terms $\varepsilon_{j}$ are non-Gaussian with strictly positive density, then the underlying graph structure is identifiable from the observed data distribution. This model is known as the linear non-Gaussian acyclic model (LiNGAM).
@@
    Hoyer et al. (2008) showed that, except in “rare” cases, a distribution does not admit an additive noise model structure in both directions simultaneously; see also Peters et al. (2017, Theorem 4.5). In fact, if the noise $\varepsilon_{Y}$ is Gaussian, then only linear functions allow for additive noise models in both directions (Hoyer et al., 2008, Corollary 1). More broadly, Peters et al. (2014) showed that if the functions $f_{j}$ in the NPSEM-IE (3) are three times differentiable and nonlinear in any component of $U_{j}$ (see Peters et al., 2014, Theorem 7.7 for a precise definition), then the corresponding graph structure is identifiable from the observed data distribution. These results extend to the so-called post-nonlinear models of the form $V_{j}=g_{j}(f_{j}(U_{j})+\varepsilon_{j})$. These models also generally exist in at most one direction, except in rare cases (Zhang and Hyvärinen, 2009).
Identification of the causal structure in linear models with Gaussian additive noise is more challenging. For example, consider again the bivariate causal model in (21). If $f_{Y}(X)$ is linear in $X$, then there exist $\beta$ and $\varepsilon_{X}$ such that $X=\beta Y+\varepsilon_{X}$ and $Y\perp\!\!\!\perp\varepsilon_{X}$ if and only if both $\varepsilon_{Y}$ and $X$ are Gaussian (Peters et al., 2017, Theorem 4.2). More generally, identification of the causal graph under linear models with Gaussian additive noise is possible if the error terms $\varepsilon_{j}$ are i.i.d., that is, the noise variances are constant across $j$ (Peters and Bühlmann, 2014). However, the equal variance assumption is not robust to rescaling of the variables.
We conclude this example with a caveat: in the nonlinear case, additive noise models are generally not closed under marginalization, even when marginalizing over unobserved variables. For instance, if $Y=f_{Y}(X)+\varepsilon_{Y}$ and $Z=f_{Z}(Y)+\varepsilon_{Z}$, then the marginal distribution $P_{X,Z}$ does not necessarily admit an additive noise model from $X$ to $Z$. This limitation restricts the applicability of additive noise models, as in most, if not all, scenarios there are intermediate variables on a causal path that are unobserved (Peters et al., 2017, §7.1.2). Therefore, additive models should perhaps be viewed as approximations to the true data-generating process rather than exact representations.
@@ 5 Summary and Practical Recommendations
The NPSEM framework formalizes causal systems using deterministic functional relationships and exogenous error terms. This allows for a generative view of how outcomes arise from their causes, enabling researchers to articulate and evaluate identification assumptions in terms of underlying data-generating mechanisms. As emphasized by Pearl (2009, p. 244), structural models and their associated graphs are particularly useful as tools for expressing assumptions about cause-effect relationships, as they provide a natural and intuitive representation of mechanistic knowledge. Counterfactual independence assumptions that may appear opaque or difficult to interpret when stated in the potential outcomes notation, such as the ignorability assumption, can often be given immediate and process-based interpretations when expressed structurally. The NPSEM framework also facilitates encoding assumptions on the functional form or even parameter values of the structural functions. However, when paired with the assumption of independent errors (NPSEM-IE), this framework often imposes strong and untestable cross-world assumptions that go beyond what is needed for identification and remain untestable even under randomized treatment assignment. These added assumptions can constrain the space of admissible models and lead to overly rigid inferences.
@@
In this workflow, structural assumptions are naturally expressed in graphical or NPSEM terms, then translated into potential outcomes notation, and finally subjected to algebraic derivation (Pearl, 2009, p. 245) to isolate those that are strictly necessary for identification. In this way, NPSEMs and their associated graphs (see Section 3.2 for details) serve as tools for evaluating the plausibility or implausibility of conditional independence assumptions stated in terms of potential outcomes. Contemporary implementations such as the target trial emulation framework (Hernán and Robins, 2016; Hernán et al., 2022) illustrate how structured causal roadmaps can help prevent common inferential pitfalls in observational studies. We refer readers to Wang et al. (2017b); Wang and Tchetgen Tchetgen (2018); Yang et al. (2019); Zhou et al. (2024); Dong et al. (2025) for examples that follow this workflow.
We now illustrate these steps in detail using Wang and Tchetgen Tchetgen (2018). Their study begins by defining the causal estimand, the ACE, in potential outcomes notation, thereby clarifying the scientific target of inference (Step 1). They then express structural assumptions through a DAG and a corresponding SWIG with bidirected arrows, explicitly distinguishing identification assumptions such as the exclusion restriction and independence from modeling assumptions (Step 2). Using algebraic derivations, they establish identification of the ACE under new no-interaction assumptions (A5.a or A5.b), showing that the estimand can be represented as the average Wald estimand (Step 3). Finally, they develop a series of bounded, efficient, and multiply robust estimators that remain consistent if any one of several working models is correctly specified (Step 4). This example demonstrates how potential outcomes, graphical models, and semiparametric theory can be systematically combined into a coherent causal workflow.
Recent advances continue to expand the reach of causal inference beyond traditional domains. Active research areas include causal discovery algorithms for learning causal structure from data (e.g., Spirtes et al., 2000; Peters et al., 2017), fairness in artificial intelligence (e.g., Kusner et al., 2017), and methods for robust learning and generalization across heterogeneous environments (e.g., Peters et al., 2016). Related ideas have also influenced reinforcement learning and decision-making in artificial intelligence, where causal reasoning is increasingly recognized as essential for interpretability and transferability. A full treatment of these topics is beyond the scope of this review, but they exemplify how causal thinking continues to shape contemporary research at the intersection of statistics, computer science, and the social and health sciences.